Calcium oxide reacts with water in a combination reaction to produce calcium hydroxide:


CaO (s) + H2O (l) → Ca(OH)2 (s)


A 3.50-g sample of CaO is reacted with 3.38 g of H2O. How many grams of water remain after completion of reaction?
1
Expert's answer
2020-04-13T02:33:30-0400

n(CaO) = "{\frac {3.50} {56}}=0.0625" mol

for reaction with this amount of CaO the same amount of H2O needed

n(H2O) reacted = 0.0625 mol

m(H2O) reacted = 0.0625*18 = 1.125 g

m(H2O) remained = 3.38 - 1.125 = 2.255 g